OpenStudy (anonymous):

In base four add the numbers 20 and 13

OpenStudy (campbell_st):

so base 4 uses digits 0, 1,2 ,3 and powers of 4 \[20 = 1 \times 4^2 +1 \times 4^1 + 0 \times 4^0\] so in when 20 is changed to base 4 it is 110

OpenStudy (campbell_st):

same idea for 13... I'll let you try

OpenStudy (campbell_st):

then when you add the 2 numbers... you carry 1 for a number larger than 4

OpenStudy (anonymous):

31

OpenStudy (campbell_st):

yep... so now you need to add